### v1

# Large Language Models with Task-Oriented Experts

Abstract
Construction of a general-purpose post-recognition error corrector poses a crucial question: how can we most effectively train a model on a large mixture of domain datasets? The answer would lie in learning dataset-specific features and digesting their knowledge in a single model. Previous methods achieve this by having separate correction language models, resulting in a significant increase in parameters. In this work, we present Mixture-of-Experts as a solution, highlighting that MoEs are much more than a scalability tool. We propose a Multi-Task Correction MoE, where we train the experts to become an “expert” of speech-to-text, language-to-text and vision-to-text datasets by learning to route each dataset’s tokens to its mapped expert. Experiments on the Open ASR Leaderboard show that we explore a new state-of-the-art performance by achieving an average relative $5.0$% WER reduction and substantial improvements in BLEU scores for speech and translation tasks. On zero-shot evaluation, NeKo outperforms GPT-3.5 and Claude-Opus with $15.5$% to $27.6$% relative WER reduction in the Hyporadise benchmark. NeKo performs competitively on grammar and post-OCR correction as a multi-task model.
[Uncaptioned image]NeKo: Toward Post Recognition Generative Correction
Large Language Models with Task-Oriented Experts
Yen-Ting Linthanks: Work done at NVIDIA research as an intern. Chao-Han Huck Yang  Zhehuai Chen Piotr Zelasko Xuesong Yang Zih-Ching Chen Krishna C Puvvada Szu-Wei Fu  Ke Hu  Jun Wei Chiu  Jagadeesh Balam  Boris Ginsburg  Yu-Chiang Frank Wang NVIDIA corresponding authors: ytl@ieee.org, hucky@nvidia.com
1 Introduction
Figure 1: Proposed NeKo, a new form multi-task model to boost post-recognition results over speech, text, and visual inputs. NeKo could work for (i) post automatic speech recognition (ASR) correction, (ii) post speech translation (ST) and machine translation (MT) correction, and (iii) post optical character recognition (OCR) correction. NeKo discover new state-of-the-art results in (iv) zero-shot ASR correction and performs competitively as a general-purpose (v) multi-task corrector.
Human recognition (Biederman, 1987; Juang and Furui, 2000; Kanwisher et al., 1996) capabilities span multiple modalities, including speech recognition, visual patterns, and extensions to semantic and textual interpretations. These faculties, however, are not infallible and often incorporate mis-recognition errors. Despite these imperfections, humans efficiently communicate using speech, language, or facial expressions.
For instance, two non-native speakers (Lev-Ari, 2015; Valaki et al., 2004) can often achieve mutual understanding through this imperfect recognition and subsequent interpretative processes, even when the conversation is marred by lexical inaccuracies and subdued accents. In other words, humans (as intelligent agents) exhibit a robust capacity for generative understanding Jiang et al. (2020); Cheng et al. (2021) that extends beyond initial recognition results.
In neuroscience Zatorre and Gandour (2008), the inferior temporal gyrus and the temporal lobe are not confined to rudimentary perception but are also integral to the post-recognition processes that facilitate semantic understanding of language (Levinson and Evans, 2010), speech (Marshall et al., 2015), and visual patterns (Vink et al., 2020). This form of “post-recognition correction,” exemplified by the application of language modeling (LM) to initial recognition outputs, has been introduced to the field for both acoustic (automatic speech recognition, ASR) and visual (optical character recognition, OCR) modalities since the early explorations (Jelinek, 1976; Dixon and Silverman, 1975) of learning algorithms in 1970s. The most prevalent approaches to utilizing LMs for post-recognition boosting are predominantly ranking or retrieval-based. In these setups, the LM is tasked with ranking and scoring (Ljolje et al., 1999) the top $n$-best hypotheses generated by the first-pass recognition system. This process often incorporates “discriminative modeling” algorithms (Sukkar and Lee, 1996; Mangu et al., 2000) and representation embeddings, such as BERT (Salazar et al., 2020; Kenton and Toutanova, 2019), to minimize the word error rate (WER) during training (Prabhavalkar et al., 2018; Mangu et al., 2000).
With the LMs scaling up to LLMs (Brown et al., 2020), recent efforts (Chan et al., 2023; Yang et al., 2023; CHEN et al., 2023; Hu et al., 2024a) have focused on exploring a “generative modeling” for post-recognition correction. This generative error correction (GER) approach uses LLMs to conduct final recognition from given first-pass text-based predictions from recognition models, including ASR, image captioning (IC), and machine translation (MT). This cascaded two-agents text-to-text GER model has outperformed larger single multi-modal and multi-task models in these tasks.
Meanwhile, these GER solutions heavily depends on domain-specific fine-tuning processes (Chen et al., 2024a) that utilize parameter-efficient components, which often suffers a performance degradation from a lack of generalizability across different datasets, domains, and tasks. In other words, how to design and further openly provide a “general-purpose post (every) recognition correction model” is still one undiscovered and crucial topic within the research community.
On the other hand, directly fine-tuning LLMs on a mixture of diverse error correction datasets can lead to suboptimal performance (CHEN et al., 2023; Lange et al., 2022) due to differences in input modalities, output formats, error types, and domain characteristics. For example, ASR errors stem from phonetic similarities or acoustic ambiguities, while OCR errors involve visual or character-level confusions. Additionally, error distributions can vary widely across datasets, even within the same task.
To characterize “model generalization,” mixture-of-experts (MoE) (Jiang et al., 2024a) has emerged as a promising approach for multi-task learning, consisting of of a set of expert networks and a gating network that learns to route the input to the most appropriate expert(Sukhbaatar et al., 2024). This enables MoE models to learn more specialized and fine-grained representations compared to monolithic models. However, most MoE models are designed for general-purpose language modeling(Dai et al., 2024), with experts not explicitly assigned to specific tasks, but rather learn to specialize in different aspects of the input space through data-driven training, expect for a recent vision work (Ye and Xu, 2023). Effectively leverage MoE for multi-task error correction, where the experts need to capture task-specific features while allowing knowledge sharing, remains an open question.
In this work, we propose NeKo, a “geNerative multi-tasK error correction” approach that leverages a pre-trained MoE model to drive diverse tasks and cross-domain knowledge, as shown in Figure 1. The key idea is to continuously pre-train MoE model on a mixture of error correction datasets, with each expert specializing in a specific domain. This task-oriented MoE fine-tuning approach enables the experts to capture task-specific features while allowing knowledge sharing through the router (Dai et al., 2024).
NeKo captures the nuances of each task, benefiting from shared knowledge across experts. Evaluated on tasks such as ASR, ST, OCR, and unseen textual error correction (TEC), NeKo consistently outperforms baseline models, including Claude-Opus and GPT-3.5. It achieves state-of-the-art WER reduction on the Hyporadise benchmark and large-scale Open ASR Leaderboard (Srivastav et al., 2023). NeKo also significant improves in OCR error correction. Further analysis confirms its robust multi-task capabilities.
In summary, the main contributions of this work include:
  1. 1.
    We introduce NeKo, a multi-task error correction LLM that leverages task-oriented mixture-of-experts for diverse post-recognition correction tasks. To the best of our knowledge, this is the first work that explores the use of MoE for multi-task error correction.
  2. 2.
    NeKo has been studied under a new form of cross-modalities post-recognition correction evaluation, serving as strong open-source ASR, ST, OCR, and TEC baselines. Our results show that NeKo discovers new state-of-the-art performance in ASR as a multi-task error correction model.
  3. 3.
    We discovered emergent abilities for cross-task correction from NeKo as a first-of-its-kind multi-task correction approach toward a general-purpose post-recognition LM designs.
  4. 4.
    The NeKo models, newly created source datasets, and training processes are scheduled to open source under the CC BY-SA 4.0 license to support reproducibility and to encourage future research.
2 Related Work
Language Modeling and Generative Error Correction
Neural correction LMs have been widely used for end-to-end (E2E) models for text error correction or normalization for both ASR (Jelinek, 1990; Irie et al., 2016; Zhang et al., 2019a, b; Guo et al., 2019) and OCR (Plamondon and Srihari, 2000; Sabir et al., 2017; Fogel et al., 2020). These models often use beam search to generate new estimates, and can usually handle text normalization and denormalization or spelling errors. To utilize the abundant textual data to improve ASR, neural LMs are integrated to E2E models using shallow fusion or rescoring over acoustic model’ likelihood confidence (Kannan et al., 2018; Salazar et al., 2020; Yang et al., 2021).
In addition to using only textual hypotheses for correction, deliberation models (Hu et al., 2020; Wang et al., 2022; Hu et al., 2023a) utilizes both audio and text for generating new ASR hypotheses. While the aforementioned methods have drawn much attention in the past, recent advances focus on using pretrained textual LLMs to benefit E2E ASR (Song et al., 2023; Hu et al., 2023c), or similar to deliberation, employing both speech and text for speech understanding by prompting (Gong et al., 2023) and joint speech and language foundation models (SLMs) (Wang et al., 2023). On the other hand, text-to-text based GEC (CHEN et al., 2023; Yang et al., 2023; Radhakrishnan et al., 2023; Chan et al., 2023) have been recently introduced to ASR, ST, image captioning, and video summarization to set up superior performance  [^1] GEC have shown better results compared to a similar scale of (i) multi-modal injection model and (ii) re-ranking based model, where we will leave these studies in the future upon having one multi-task correction LM as NeKo presented here. Please also. These GEC approaches take the text output as an input to large language models (LLMs), enabling them to perform zero-shot typo correction by generatively refining the final recognition results (CHEN et al., 2023; Yang et al., 2023; Radhakrishnan et al., 2023; Chan et al., 2023). In this work, we aim to design a text-to-text model as a first step for providing open-source post-recognition correction to the communities.
Mixture of Experts
Mixture-of-experts (MoE) (Shazeer et al., 2017) is a machine learning concept that employs multiple expert layers, each of which specializes in solving a specific subtask. The experts then work together to solve the entire task at hand. Recently, MoE has been widely applied to large-scale distributed Deep Learning models by using a cross-GPU layer that exchanges hidden features from different GPUs (Lepikhin et al., 2021; Fedus et al., 2022). The MoE approach is differentiated from existing scale-up approaches for DNNs, such as increasing the depth or width of DNNs, in terms of its high cost-efficiency. Specifically, adding more model parameters (experts) in MoE layers does not increase the computational cost per token at inference time. Thus, MoE has been studied for scaling the models to trillion-size parameters in NLP (Fedus et al., 2022). In speech processing, Gaur et al. (2021); Hu et al. (2023b); You et al. (2021) primarily showed benefits of MoE on speech recognition tasks. In this work, we further pursue this direction by modeling MoE on error correction and highlight the effectiveness and robustness of MoEs in learning from a mixture of correction datasets.
3 Method
Figure 2: The architecture of our proposed model, NeKo, which integrates MoE layers within a Transformer architecture. During inference, we do not assume knowledge of the specific task an input belongs to and each token is routed to the top-$2$ experts solely based on their router probabilities.
3.1 Mixture-of-Experts (MoE)
Our method, NeKo, is based on a Transformer architecture (Vaswani et al., 2017) with modifications similar to those described in Jiang et al. (2023). The key difference is that we replace the feedforward blocks with Mixture-of-Expert (MoE) layers. In a MoE layer, each input token is assigned to a subset of experts by a gating network (router). The output of the MoE layer is the weighted sum of the outputs of the selected experts, where the weights are determined by the gating network. Formally, given $n$ expert networks $\{E_{0},E_{1},...,E_{n-1}\}$, the output of the MoE layer for an input token $x$ is:
$y=\sum_{i=0}^{n-1}G(x)_{i}\cdot E_{i}(x),$ (1)
where $G(x)_{i}$ is the weight assigned to the $i$-th expert by the gating network, and $E_{i}(x)$ is the output of the $i$-th expert network for input $x$. The gating network $G(x)$ is implemented as a softmax over the top-$K$ logits of a linear layer:
$G(x)=\text{Softmax}(\text{TopK}(x\cdot W_{g})),$ (2)
where $\text{TopK}(\ell)_{i}=\ell_{i}$ if $\ell_{i}$ is among the top-$K$ coordinates of logits $\ell\in\mathbb{R}^{n}$, and $\text{TopK}(\ell)_{i}=-\infty$ otherwise. The number of experts $K$ used per token is a hyperparameter that controls the computational cost.
3.2 Task-Oriented Expert Assignment
The key idea of NeKo is to assign each expert to a specific task during training. Given a set of tasks $\mathcal{T}=\{T_{1},T_{2},...,T_{m}\}$, we define a mapping function $f:\mathcal{T}\rightarrow\{1,2,...,n\}$ that assigns each task to a unique expert. During training, for an input token $x$ from task $T_{i}$, we deterministically route $x$ to the expert $f(T_{i})$ in addition to the top-1 expert selected by the gating network. This ensures that each expert learns task-specific features while still allowing for knowledge sharing through the gating network. Formally, the output of the MoE layer for an input token $x$ from task $T_{i}$ during training is:
$y=G(x)_{f(T_{i})}\cdot E_{f(T_{i})}(x)+G(x)_{\text{top1}}\cdot E_{\text{top1}}(x),$ (3)
[…]
During inference, we do not assume knowledge of the specific task an input token belongs to. Instead, we route each token to the top-$K$ experts selected by the gating network based on their predicted probabilities. This approach allows the model to leverage the task-specific knowledge learned by the experts during training while still being able to generalize to new, potentially unseen tasks and domains during inference.
3.3 Training Objective
We train NeKo on a mixture of error correction datasets $\mathcal{D}=\{D_{1},D_{2},...,D_{m}\}$, where each dataset $D_{i}$ corresponds to a specific task $T_{i}$. The training objective is to minimize the negative log-likelihood of the target sequences:
$\mathcal{L}=-\sum_{i=1}^{m}\sum_{(x,y)\in D_{i}}\log p(y|x,T_{i}),$ (4)
where $x$ is the input sequence (e.g., ASR hypotheses, OCR output), $y$ is the target sequence (e.g., ground-truth transcription, corrected text), and $p(y|x,T_{i})$ is the probability of the target sequence given the input sequence and the task prompt (Figure 3.) By jointly training on multiple error correction datasets with task-oriented expert assignment, NeKo learns to capture task-specific features while allowing for knowledge sharing across tasks through the shared gating network and other model components.
4 Experiments
4.1 Training and Evaluation Datasets
ASR
To assess the ability to handle diverse and noisy real-world speech, we use the Open ASR Leaderboard (Gandhi et al., 2022; Srivastav et al., 2023) for ASR evaluation, which comprises nine diverse datasets spanning various domains and speaking styles. These include LibriSpeech (Panayotov et al., 2015), Common Voice 9 (Ardila et al., 2020), VoxPopuli (Wang et al., 2021), TED-LIUM (Hernandez et al., 2018), GigaSpeech (Chen et al., 2021), SPGISpeech (O’Neill et al., 2021), Earnings-22 (Del Rio et al., 2022), and AMI (Carletta, 2007; Renals et al., 2007), as one most representative benchmark due to its scale and data diversity. We include the training set of above 8 datasets for NeKo training. We use the word error rate as the evaluation metric for ASR.
ST and MT
For the speech translation error correction task, we use the subset of the HypoTranslate dataset (Hu et al., 2024b) for training and evaluation. This dataset includes translation results from FLEURS (Conneau et al., 2022), CoVoST-2 (Wang et al., 2020), and MuST-C (Di Gangi et al., 2019), covering a range of languages such as Spanish, French, Italian, Japanese, Portuguese, Chinese, and Persian. As an extra zero-shot textual correction setup, we evaluate NeKo on machine translation (MT) of WMT’20 for Japanese and Chinese (Barrault et al., 2020a). We use the BLEU score (Papineni et al., 2002) as the evaluation metric for ST (with training and test) and MT (zero-shot).
OCR
For the optical character recognition (OCR) error correction task, we use the English portion of the Post-OCR Correction dataset (PleIAs, 2023), which contains newspaper texts from Chronicling America. The dataset includes original texts with varying numbers of OCR mistakes and their corresponding corrected versions. To evaluate our model, we take the first 1,000 characters of both the input text with OCR errors and the ground-truth corrected text. We use the WER as the evaluation metric.
TEC
For the textual error correction (TEC) task, we use a subset of the CoEdIT dataset (Raheja et al., 2023), which contains a diverse collection of 82K task-specific instructions for text editing. We select two editing tasks from CoEdIT that align with our error correction objectives: grammar correction and coherence improvement. These tasks focus on correcting grammatical errors and improving the overall coherence of the text, making them suitable for evaluating the effectiveness of our model in handling TEC-related editing instructions. We use the word error rate as the evaluation metric.
4.2 Task-Specific Recognition Systems and Baselines
ASR
We compare against state-of-the-art ASR models, Distil-Whisper-V2-Large(Gandhi et al., 2023), Whisper-V2-Large, Whisper-V3-Large(Radford et al., 2022), Canary(NVIDIA, 2024) without applying GEC method. A end-to-end ASR-LLM, SALM (Chen et al., 2024b), is also compared. For all Cascaded ASR+GEC Methods, the task-specific system is the Canary model. This model transcribes the speech data and generate 5-best hypotheses for each utterance using temperature-based sampling (Ackley et al., 1985) with $p=0.3$. This allows us to capture a diverse set of potential transcriptions for each utterance, which can then be fed into our error correction model.
ST and MT
For the speech and machine translation tasks, we compare against state-of-the-art models SeamlessM4T (Barrault et al., 2023a), GenTranslate(Hu et al., 2024c), and cascaded approaches combining ASR and machine translation models (e.g., Whisper + NLLB (Costa-jussà et al., 2022)). These baselines cover both end-to-end speech translation models and pipeline approaches. We use SeamlessM4T-Large V2 as the task-specific system to decode $N$-best hypotheses from input speech by beam search algorithm. We did this in two steps by first transcribing the speech and then translating the text, following (Hu et al., 2024c). LLMs then take the N-best hypotheses to produce a final speech translation result. To investigate the generalization of our model, we also evaluate it in an alternative scenario: a direct speech translation model, Canary, is used as the task-specific system to produce hypotheses.
OCR and TEC
We compare our proposed method against two baselines: (1) the input text without any correction (denoted as Baseline) and (2) a Mixtral 8x7B model fine-tuned only on the respective dataset for each task (denoted as Mistral 8x7B Direct Finetune). This allows us to assess the effectiveness of our task-oriented expert assignment approach in handling OCR and TEC errors, as well as its ability to leverage knowledge from multiple tasks to improve performance on individual tasks compared to direct fine-tuning on a single dataset.
Table 1: Cross-domain ASR correction results in zero-shot and few-shot settings on the Hyporadise benchmark (CHEN et al., 2023). We compare NeKo against GPT-3.5 Turbo and Claude-Opus in 0- and 5-shot settings. The baseline represents the WER of task-specific model Whisper-Large. The oracle results used in CHEN et al. (2023) (N-best and Compositional) provide an upper bound for the correction performance.
Domain Test Set Baseline GPT-3.5 Turbo Claude-Opus 0-shot w/ NeKo Oracle
Shift 0-shot 5-shot 0-shot 5-shot NeKo-FFT NeKo-BTX NeKo-MoE N-best Comp.
Specific Scenario WSJ-dev93 9.0 $8.5_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $7.7_{{\color[rgb]{0,.5,.5}-14.4\%}}$ $8.2_{{\color[rgb]{0,.5,.5}-8.9\%}}$ $7.4_{{\color[rgb]{0,.5,.5}-17.8\%}}$ $8.6_{{\color[rgb]{0,.5,.5}-4.4\%}}$ $7.5_{{\color[rgb]{0,.5,.5}-16.7\%}}$ $\textbf{6.8}_{{\color[rgb]{0,.5,.5}-24.4\%}}$ 6.5 5.3
WSJ-eval92 7.6 $7.3_{{\color[rgb]{0,.5,.5}-3.9\%}}$ $6.6_{{\color[rgb]{0,.5,.5}-13.2\%}}$ $7.0_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $6.3_{{\color[rgb]{0,.5,.5}-17.1\%}}$ $7.4_{{\color[rgb]{0,.5,.5}-2.6\%}}$ $6.4_{{\color[rgb]{0,.5,.5}-15.8\%}}$ $\textbf{5.8}_{{\color[rgb]{0,.5,.5}-23.7\%}}$ 5.5 4.7
ATIS 5.8 $5.5_{{\color[rgb]{0,.5,.5}-5.2\%}}$ $5.0_{{\color[rgb]{0,.5,.5}-13.8\%}}$ $5.2_{{\color[rgb]{0,.5,.5}-10.3\%}}$ $4.7_{{\color[rgb]{0,.5,.5}-19.0\%}}$ $5.6_{{\color[rgb]{0,.5,.5}-3.4\%}}$ $4.8_{{\color[rgb]{0,.5,.5}-17.2\%}}$ $\textbf{4.2}_{{\color[rgb]{0,.5,.5}-27.6\%}}$ 3.5 2.4
Common Noise CHiME4-bus 18.8 $17.6_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $16.2_{{\color[rgb]{0,.5,.5}-13.8\%}}$ $17.1_{{\color[rgb]{0,.5,.5}-9.0\%}}$ $15.7_{{\color[rgb]{0,.5,.5}-16.5\%}}$ $17.7_{{\color[rgb]{0,.5,.5}-5.9\%}}$ $15.9_{{\color[rgb]{0,.5,.5}-15.4\%}}$ $\textbf{14.5}_{{\color[rgb]{0,.5,.5}-22.9\%}}$ 16.8 10.7
CHiME4-caf 16.1 $14.7_{{\color[rgb]{0,.5,.5}-8.7\%}}$ $13.7_{{\color[rgb]{0,.5,.5}-14.9\%}}$ $14.2_{{\color[rgb]{0,.5,.5}-11.8\%}}$ $13.2_{{\color[rgb]{0,.5,.5}-18.0\%}}$ $14.8_{{\color[rgb]{0,.5,.5}-8.1\%}}$ $13.4_{{\color[rgb]{0,.5,.5}-16.8\%}}$ $\textbf{12.2}_{{\color[rgb]{0,.5,.5}-24.2\%}}$ 13.3 9.1
CHiME4-ped 11.5 $10.9_{{\color[rgb]{0,.5,.5}-5.2\%}}$ $9.7_{{\color[rgb]{0,.5,.5}-15.7\%}}$ $10.5_{{\color[rgb]{0,.5,.5}-8.7\%}}$ $9.3_{{\color[rgb]{0,.5,.5}-19.1\%}}$ $11.0_{{\color[rgb]{0,.5,.5}-4.3\%}}$ $9.5_{{\color[rgb]{0,.5,.5}-17.4\%}}$ $\textbf{8.6}_{{\color[rgb]{0,.5,.5}-25.2\%}}$ 8.5 5.5
CHiME4-str 11.4 $10.9_{{\color[rgb]{0,.5,.5}-4.4\%}}$ $9.7_{{\color[rgb]{0,.5,.5}-14.9\%}}$ $10.5_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $9.3_{{\color[rgb]{0,.5,.5}-18.4\%}}$ $11.0_{{\color[rgb]{0,.5,.5}-3.5\%}}$ $9.4_{{\color[rgb]{0,.5,.5}-17.5\%}}$ $\textbf{8.5}_{{\color[rgb]{0,.5,.5}-25.4\%}}$ 9.0 6.0
Speaker Accent MCV-af 25.3 $24.9_{{\color[rgb]{0,.5,.5}-1.6\%}}$ $23.6_{{\color[rgb]{0,.5,.5}-6.7\%}}$ $24.4_{{\color[rgb]{0,.5,.5}-3.6\%}}$ $23.0_{{\color[rgb]{0,.5,.5}-9.1\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-1.2\%}}$ $23.3_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $\textbf{21.0}_{{\color[rgb]{0,.5,.5}-17.0\%}}$ 23.6 21.7
MCV-au 25.8 $25.1_{{\color[rgb]{0,.5,.5}-2.7\%}}$ $24.0_{{\color[rgb]{0,.5,.5}-7.0\%}}$ $24.6_{{\color[rgb]{0,.5,.5}-4.7\%}}$ $23.4_{{\color[rgb]{0,.5,.5}-9.3\%}}$ $25.2_{{\color[rgb]{0,.5,.5}-2.3\%}}$ $23.7_{{\color[rgb]{0,.5,.5}-8.1\%}}$ $\textbf{21.4}_{{\color[rgb]{0,.5,.5}-17.1\%}}$ 24.9 21.8
MCV-in 28.6 $27.6_{{\color[rgb]{0,.5,.5}-3.5\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-12.6\%}}$ $27.0_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $24.3_{{\color[rgb]{0,.5,.5}-15.0\%}}$ $27.8_{{\color[rgb]{0,.5,.5}-2.8\%}}$ $24.6_{{\color[rgb]{0,.5,.5}-14.0\%}}$ $\textbf{22.2}_{{\color[rgb]{0,.5,.5}-22.4\%}}$ 27.1 22.6
MCV-sg 26.4 $26.5_{{\color[rgb]{.5,.5,.5}+0.4\%}}$ $25.1_{{\color[rgb]{0,.5,.5}-4.9\%}}$ $25.9_{{\color[rgb]{0,.5,.5}-1.9\%}}$ $24.5_{{\color[rgb]{0,.5,.5}-7.2\%}}$ $26.6_{{\color[rgb]{.5,.5,.5}+0.8\%}}$ $24.7_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $\textbf{22.3}_{{\color[rgb]{0,.5,.5}-15.5\%}}$ 25.5 22.2
Table 2: ASR correction results on the Open ASR Leaderboard. We report the Word Error Rate (WER) for each dataset and the average across all 9 datasets. NeKo establishes a new state-of-the-art performance on the leaderboard, outperforming both end-to-end ASR methods and cascaded ASR+GEC approaches. Lower WER indicates better performance.
Model AMI Earnings22 Gigaspeech LS Clean LS Other SPGI Tedlium Voxp. MCV9 Average
End-to-end ASR Methods
Distil-Whisper-V2-Large (Gandhi et al., 2023) 14.65 12.12 10.31 2.95 6.39 3.28 4.30 8.22 12.60 8.31
Whisper-V2-Large (Radford et al., 2022) 16.82 12.02 10.57 2.56 5.16 3.77 4.01 7.50 10.11 8.06
Whisper-V3-Large 16.01 11.30 10.02 2.03 3.91 2.95 3.90 9.52 9.67 7.70
Canary (NVIDIA, 2024) 14.00 12.25 10.19 1.49 2.49 2.06 3.58 5.81 7.75 6.67
Bestow Speech LM (Chen et al., 2024c) 12.58 12.86 10.06 1.64 3.07 2.11 3.41 5.84 6.97 6.50
Cascaded ASR+GEC Methods
Gemma 2B (Team et al., 2024) FFT 13.20 12.30 10.40 1.60 2.60 2.20 3.70 6.00 7.50 6.61
Gemma 8x2B FFT 13.10 12.20 10.30 1.50 2.50 2.10 3.60 5.90 7.40 6.51
NeKo (Ours) Gemma 8x2B 13.00 12.10 10.20 1.40 2.40 2.00 3.50 5.80 7.30 6.41
Mistral 7B (Jiang et al., 2023) FFT 13.07 11.87 10.09 1.48 2.46 2.04 3.55 5.75 7.29 6.40
Mixtral 8x7B (Jiang et al., 2024b) FFT 12.91 12.19 10.34 1.54 2.55 2.12 3.64 5.89 7.43 6.51
Mixtral 8x7B Lora 12.96 12.24 10.38 1.55 2.56 2.13 3.66 5.92 7.47 6.60
Mistral 8x7B Branch-Train-Mix (Sukhbaatar et al., 2024) 13.13 11.93 10.14 1.49 2.47 2.05 3.57 5.78 7.33 6.43
NeKo (Ours) Mixtral 8x7B 12.55 11.82 10.02 1.49 2.47 2.05 3.52 5.76 7.25 6.34
NeKo (Ours) Mixtral 8x22B 12.61 11.93 10.15 1.52 2.51 2.09 3.58 5.82 7.33 6.40
4.3 Post-recognition LLMs Setup
We implement NeKo using the Transformer architecture (Vaswani et al., 2017) and fine-tune both dense and MoE models for comparison. For dense models, we fine-tune Gemma 2B (Team et al., 2024) and Mistral 7B (Jiang et al., 2024b). For MoE models, we fine-tune Gemma 8x2B [^2] We made an up-cycled (Komatsuzaki et al., 2023) Gemma 8x2B MoE setup extended from single Gemma-2B (Team et al., 2024). and Mixtral 8x7B without applying our task-oriented expert assignment. We explore the Branch-Train-Mix approach (Sukhbaatar et al., 2024), which involves branching from the Mistral 7B model to an 8x7B MoE model as one competing setup. To investigate the scalability of our method, we design NeKo to three different sizes of MoE models: Gemma 8x2B, Mixtral 8x7B, and Mixtral 8x22B. We further compared low-rank adaptation (LoRA(Hu et al., 2021)) with full fine-tuning (FFT) on 8x7B MoE setup.
For MoE models, we use top-k routing as proposed in (Lepikhin et al., 2021) to balance the computational cost and model capacity. We use a global batch size of 2 million tokens and apply sample packing (Raffel et al., 2020) to maximize the GPU utilization. We fine-tune the model for 3 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $41e-$ 4 and a weight decay of $0.01$. We use a cosine learning rate scheduler with a warmup ratio of $0.1$ and a gradient clipping threshold of $1.0$. For the expert-dataset mapping, we randomly assign each dataset to one of the 8 experts in the Mixtral model. This random assignment serves as a strong baseline and allows us to focus on the effectiveness of the task-oriented expert assignment approach. We leave the exploration of more advanced expert assignment strategies for future work. To efficiently train the large-scale model, we leverage DeepSpeed Zero (Rajbhandari et al., 2020) for memory optimization and Hugging Face Transformers (Wolf et al., 2020) for model implementation.
[…]
Es Fr It Ja Pt Zh Avg. Fa Ja Zh Avg. Es It Zh Avg.
End-to-end ST Methods
SeamlessM4T-Large (Barrault et al., 2023a) 23.8 41.6 23.9 21.0 40.8 28.6 30.0 18.3 24.0 34.1 25.5 34.2 29.9 16.2 26.8
GenTranslate (Hu et al., 2024c) 25.4 43.1 25.5 28.3 42.4 34.3 33.2 21.1 29.1 42.8 31.0 33.9 29.4 18.5 27.3
SeamlessM4T-Large-V2 (Barrault et al., 2023b) 23.8 42.6 24.5 21.7 43.0 29.5 30.9 16.9 23.5 34.6 25.0 32.1 27.5 15.6 25.1
GenTranslate-V2 (Hu et al., 2024c) 25.5 44.0 26.3 28.9 44.5 34.9 34.0 19.4 29.0 43.6 30.7 32.2 27.3 18.1 25.9
Cascaded ASR+MT Methods
Whisper + NLLB-3.3b (Costa-jussà et al., 2022) 25.1 41.3 25.0 19.0 41.5 23.5 29.2 13.6 19.0 32.0 21.5 35.3 29.9 13.5 26.2
SeamlessM4T-Large (ASR+MT) (Barrault et al., 2023a) 24.6 44.6 25.4 22.5 41.9 31.2 31.7 18.8 24.0 35.1 26.0 35.1 30.8 17.7 27.9
SeamlessM4T-V2 (ASR+MT) (Barrault et al., 2023b) 24.7 44.1 25.1 20.6 43.6 30.6 31.5 17.4 23.8 35.4 25.5 33.0 27.8 14.5 25.1
Cascaded ASR+GEC Methods
GenTranslate 26.8 45.0 26.6 29.4 43.1 36.8 34.6 21.8 30.5 43.3 31.9 35.5 31.0 19.6 28.7
GenTranslate-V2 27.0 44.3 26.4 27.8 44.5 36.1 34.4 20.8 29.7 43.5 31.3 33.2 28.3 16.9 26.1
NeKo-FT 26.9 44.2 26.3 27.7 44.4 36.0 34.3 20.7 29.6 43.4 31.2 33.1 28.2 16.8 26.0
NeKo-BTX 27.2 44.5 26.7 28.0 44.7 36.3 34.6 21.0 29.9 43.8 31.6 33.4 28.5 17.1 26.3
NeKo-MoE 28.5 46.2 28.0 30.1 46.3 38.7 36.3 23.4 32.6 46.5 34.2 37.2 32.8 21.5 30.5
4.4 Post-recognition Generative Correction Results
ASR
We first evaluate the zero-shot ability of NeKo on unseen domain compared to two general-purpose LLMs, including GPT-3.5 Turbo and Claude-Opus [^3] Claude-Opus is one of the high-performance, general-purpose LLMs evaluated in the LLM evaluation (Zheng et al., 2024) and performs competitively compared to GPT-4-Turbo and Gemini 1.5-Pro, offering unlimited API access.. With a task-specific recognition baseline of Whisper-V2-Large (third column) in Table 1, NeKo-MoE shows the best zero-shot ability with a relative 22.3% average WER reduction. GPT-3.5 Turbo and Claude-Opus have relative 4.3% and 7.3% of zero-shot improvements, where NeKo consistently outperform their 5-shot ASR correction.
Table 2 shows the WER scores on individual datasets and average performance on the Open ASR Leaderboard. We observe that the proposed NeKo improves the task-specific baseline Canary, with an average 5.0% WER reduction. Individually, we observe a significant performance increase with NeKo on more challenging datasets, like AMI (conversational speech) and VoxPopuli (accented speech) due to experts learning dataset-specific features. While, Earnings22 shows a slight performance drop possibly due to the reduced representation in the batch.
Compared to other leading models on the leaderboard, NeKo establishes a new state-of-the-art, outperforming speech-only foundational models like Whisper and Canary and end-to-end ASR-LLM like SALM  [^4] We conduct one competitive end-to-end, multi-modal baseline based on Bestow (Chen et al., 2024c) with joint audio and text injection under cross-attention, which shows improved performance on the Open ASR benchmark. (Chen et al., 2024b) across most datasets. On the AMI dataset, NeKo achieves a WER of 12.58%, significantly lower than Whisper’s 16.82%. On VoxPopuli, NeKo obtains 5.84% WER, a 1.66 point reduction from Whisper’s 7.5%. The strong performance of NeKo demonstrates the effectiveness of our speech-adapted MoE approach in handling diverse speech datasets and learning robust representations.
ST and MT
Table 3 presents the speech translation results on the FLEURS, CoVoST-2, and MuST-C datasets. For these experiments, we use SeamlessM4T-Large as the task-specific model to generate the initial speech translation hypotheses. NeKo is then applied to correct the outputs from SeamlessM4T-Large. Compared to the task-specific SeamlessM4T-Large model, NeKo achieves significant improvements, with an average BLEU score increase of 5.4 points on the FLEURS dataset, 9.2 points on the CoVoST-2 dataset, and 5.4 points on the MuST-C dataset. These results demonstrate the effectiveness of NeKo in correcting errors made by the first-pass speech translation model. Moreover, NeKo outperforms other correction baselines, including the state-of-the-art GenTranslate model. On the FLEURS dataset, NeKo obtains an average BLEU score of 36.3, surpassing GenTranslate by 2.3 points. For CoVoST-2, NeKo achieves an average BLEU score of 34.2, outperforming GenTranslate by 3.2 points. On MuST-C, NeKo improves over GenTranslate by 3.2 BLEU points, reaching an average score of 30.5.
Table 4: Machine translation BLEU scores on the WMT’20 Japanese (Ja) and Chinese (Zh) test sets (Barrault et al., 2020b). NeKo is evaluated in a zero-shot setting, while other models are fine-tuned on the respective language pairs. Higher BLEU scores indicate better translation quality.
[…]
ALMA-13b   3.5 11.3   7.4
BigTranslate   7.3 29.0 18.2
NLLB-3.3b 11.6 26.9 19.3
SeamlessM4T-Large 17.0 27.0 22.0
GenTranslate (fine-tuned) 21.4 30.7 26.1
NeKo-MoE (zero-shot) 18.1 27.6 22.9
To further assess the generalization ability of NeKo , we evaluate it on the WMT’20 machine translation benchmark for Japanese and Chinese in a zero-shot setting. As shown in Table 4, NeKo achieves competitive performance compared to fine-tuned MT models, obtaining an average BLEU score of 22.9. This result highlights the potential of NeKo to handle unseen translation tasks by leveraging the knowledge learned from speech translation datasets.
OCR and TEC
For the OCR task, NeKo achieves a substantial error reduction, lowering the WER from 71.03% to 14.43%. This represents a significant improvement over the baseline and demonstrates the model’s ability to correct OCR errors effectively. Compared to the Mixtral-MoE model fine-tuned directly on the OCR dataset, NeKo obtains a 1.02% lower WER, highlighting the benefit of the task-oriented expert assignment approach. In the TEC task, NeKo showcases its versatility by improving the performance on both grammar correction and coherence improvement subtasks. For grammar correction, NeKo reduces the WER from 31.41% to 9.42%, outperforming the directly fine-tuned Mixtral-MoE model by 1.31%. On the coherence subtask, NeKo achieves a WER of 9.71%, which is 0.46% higher than the directly fine-tuned model but still a significant improvement over the baseline.
Table 5: WER comparison of NeKo against the baseline and a directly fine-tuned Mixtral-MoE model (8x7B) on grammar correction and coherence improvement tasks from the CoEdIT dataset (Raheja et al., 2023), and the OCR task using the PleIAs/Post-OCR-Correction dataset (PleIAs, 2023). Lower WER indicates better performance.
Task Grammar Correction Coherence Improv. OCR
Baseline 31.41 13.48 71.03
Mixtral-MoE-FFT 10.73   9.25 15.45
NeKo-MoE   9.42   9.71 14.43
Table 6: WER comparison of NeKo against GPT-3.5-Turbo, and Claude-Opus on the 5-shot IMDb typographical error correction dataset (Shah and de Melo, 2020). The baseline represents the WER between the corrupted text and the ground truth. Lower WER indicates better performance in correcting typographical errors.
Model WER
Baseline (Corrupt vs Ground Truth) 18.35%
GPT-3.5-Turbo (5-shots) 12.72%
Claude-Opus (5-shots) 8.18%
NeKo-MoE (5-shots) 11.62%
Emergent Unseen Task Zero-Shot Performance
We investigate NeKo’s generalization capabilities to unseen tasks using an additional synthetic typographical error correction dataset (Shah and de Melo, 2020). This dataset is derived from the IMDb test split, featured low noise levels (3.75% character error rate) with corruption applied using algorithms proposed in (Shah and de Melo, 2020). Our evaluation focused on zero-shot and five-shot learning scenarios to assess the adaptability of various models without and with minimal task-specific training. In the zero-shot scenario, where models were prompted to switch from an ASR task to typo correction without additional training, the challenge proved significant. The models, including the advanced Claude-Opus, yielded WERs above 30%. The predictions were markedly irrelevant to the ground truth, highlighting the difficulty of adapting to typo correction without specific fine-tuning. This finding prompts further investigation into efficient and effective training techniques for generalizing model capabilities across diverse linguistic tasks. In the five-shot scenario, all models improved against the corrupted baseline with Claude-Opus performing best. Notably, NeKo outperformed GPT-3.5-Turbo, indicating some affinity towards this task.
5 Conclusion
In this work, we proposed NeKo, a multi-task GER approach that leverages task-oriented MoEs to handle diverse tasks. NeKo assigns each expert to a specific dataset during training, enabling the experts to capture task-specific features while allowing knowledge sharing through the gating network. Our results show that task-oriented expert assignment is a promising approach for multi-task learning in error correction and other natural language processing tasks. By aligning experts with datasets, NeKo can effectively capture the nuances and specificities of each task while benefiting from the shared knowledge learned by the gating network and other model components. Future work includes exploring more advanced expert assignment strategies, such as dynamically assigning experts based on the input characteristics. Investigating the interpretability of the learned expert representations and routing decisions would be important next step upon the open NeKo models.
Limitation
We aim to provide a transparent and comprehensive understanding of the current scope of NeKo, and pave the way for future research to further improve the NeKo model.
Dataset Diversity and Size and Assumptions in Error Distribution
This study addresses a mixture of error correction tasks, including ASR, ST, OCR, and TEC, using representative task-specific datasets such as LibriSpeech for ASR, CoVoST for ST, ICDAR 2019 for OCR, and CoNLL-2014 for TEC. While these datasets are widely recognized benchmarks, they may not cover all possible error correction scenarios, particularly those involving more complex or less common error types found in real-world data. This setup assumes that the error distributions in the training datasets are representative of those in real-world applications. Consequently, the performance of NeKomight be overestimated for certain types of data not covered by these benchmarks, affecting the generalizability of the results to more diverse and noisy real-world scenarios. Future research should include a broader range of datasets, particularly those with more diverse and challenging error types, and investigate methods to dynamically adapt to varying error distributions, possibly through online learning Yasunaga et al. (2021) or domain adaptation techniques Khurana et al. (2021), to better evaluate the robustness and generalizability of the model.
Ethical and Societal Considerations
The study does not extensively address the ethical and societal implications of deploying NeKo in real-world applications. There could be unintended consequences, such as biases in error correction or misuse of the technology in sensitive applications. Future work should include a thorough analysis of the ethical and societal impacts of the model, along with strategies to mitigate potential negative consequences. This could involve incorporating fairness and bias detection mechanisms Liu et al. (2022) into the model to ensure responsible and ethical deployment.
Boarder Impacts
The NeKo model’s application of MoE for multi-domain and multi-task error correction has the potential to significantly enhance automated system’s performance across various domains, such as healthcare, education and customer service. By improving standard mediums of communication such as speech recognition, translation and optical character recognition NeKo can facilitate more inclusive technologies, benefiting individuals with impairments or non-native speakers. Additionally, the economic benefits from reduced manual correction efforts and educational advantages from more accurate communication system can be substantial. The open-sourcing of NeKo under the CC BY-SA 4.0 license encourages collaboration and reproducibility with in the reserach community, fostering innovation and broader application. Future work should also consider optimizing the training process to minimize the environmental impact, promoting sustainable AI development practices.
Task-Specific Fine-Tuning
The NeKo model employs task-oriented MoE fine-tuning, where each expert is assigned to a specific dataset. This approach may lead to overfitting to the specific characteristics of the training datasets even though knowledge could be shared. As a result, the model’s performance might degrade when applied to new tasks or datasets that were not part of the training set, limiting its adaptability. Investigating more dynamic and adaptive fine-tuning strategies that can generalize better across unseen tasks and datasets would be beneficial. Techniques such as meta-learning or continual learning could be explored to enhance the model’s adaptability and robustness.
Future Connections to In-Context and Auto-Agent Learning with NeKo
Integrating in-context learning (ICL) with NeKo could enable the model to adapt to various error correction tasks by conditioning on input examples without requiring explicit fine-tuning. This approach is particularly beneficial in scenarios where obtaining large labeled datasets for fine-tuning is impractical. By leveraging ICL, NeKo could adapt to diverse error types and use in-context examples to correct errors specific to new domains or applications, thereby improving its generalizability to real-world data. Furthermore, ICL would allow the model to dynamically adjust its error correction strategies based on the input context, enhancing its robustness to varying error distributions.
Appendix A Appendix
Prompt Format
We provide detailed correction example per [TASK] and actual prompt format of INPUT: used in the our experiments for qualitative studies as shown in Figure 3. For instance, each task will have a specific task-activation prompt format, where ASR, ST, and MT would be based on the sampling or beam search results. On the other hand, OCR and TEC will use input texts for end-to-end mapping.
Figure 3: Example prompts of various correction tasks using Automatic Speech Recognition (ASR), Machine Translation (MT), Speech Translation (ST), Optical Character Recognition (OCR), and Textual Error Correction (TEC).
Correction Examples
We randomly select post-recognition example by NeKo . In Figure 4, a long form ASR output has been selected and it remain the top 1-best correction with NeKo. or the ST and MT correction result in Figure 5 and in Figure 6, although the post-NeKo corrected output does not perfectly align with the ground truth, it boosts the general semantic meaning, as reviewed by native speakers. Meanwhile, the OCR and TEC correction results in Figures  7 and  8 demonstrate various types of corrections, such as pattern-wise character misrecognition and understanding-based coherence improvements.
Figure 4: Examples of NeKo outputs for asr error correction task in SPGISpeech O’Neill et al. (2021).
Figure 5: Examples of NeKo outputs for speech translation correction task in FLEURS Conneau et al. (2022).
Figure 6: Examples of NeKo outputs for machine translation correction task in WMT20 Barrault et al. (2020a).
Additional Discussion on Human Recognition from Speech and Text Inputs
Human recognition (e.g., speech, optical character, text translation) and has naturally evolved to excel at recognizing and understanding speech in a wide range of real-world scenarios He et al. (2019); Deng et al. (2013). However, the field of automatic speech recognition (ASR) has traditionally concentrated on training and evaluating models on specific datasets Chan et al. (2016); Watanabe et al. (2017). These models have shown limited adaptability to new environments Yang et al. (2021); Du et al. (2016); Hu et al. (2024a), leading to decreased accuracy and practicality in real-world settings. Recognizing the challenges posed by single dataset models and the availability of diverse datasets collected over time, unified models are being developed that merge information from multiple datasets into a single framework Barrault et al. (2023a). While Grammatical Error Correction (TEC) has been actively explored Yang et al. (2023), ASR error correction is distinct due to the arbitrariness of spoken language Aksënova et al. (2021), requiring efforts from both speech, NLP, and cognitive science communities as one human recognition example shown in Figure 9.
Task-Oriented Inference for Mixture of Expert Models
During inference, the Neko-model utilizes top-2 expert routing, instead of just top-1. Our pilot studies showed that top-1 routing indeed led to worse performance due to limited knowledge sharing.
Using more than two experts (e.g., top-3 or higher) diverged from the training setup and increased inference costs (ranging from $23.5$% to $75.5$%) without significant gain (i.e., a relative difference of less than $0.06$%).
Future Model Maintenance Plan and ASR Community
For ASR tasks, we used Canary-v0, Whisper-seires, and SeamlessM4T to decode textual hypotheses data. For Whisper, we included it as a widely-used baseline, but our key comparisons are to other GEC methods also using Whisper (e.g. GenTranslate). Open eco-system, including ESPnet (Watanabe et al., 2018) and SpeechBrain (Ravanelli et al., 2021) models, are also our interests to be adapted as first-pass ASR in the open code base. This will provide a more comprehensive evaluation across model types. In general, NeKo’s post-ASR correction improvements are consistent across datasets and first-pass models, suggesting the benefits generalize beyond model-specific (i.e., Canary, Whisper, or SeamlessM4T) ’s strengths as the initial medical term correction results shown in Figure 10.
Acknowledgment
The authors would like to acknowledge the initial feedback from Taejin Park, Boyi Li, and Yejin Choi at NVIDIA and the discussion on medical speech recognition evaluation with Zhen Wan and Chenhui Chu from Kyoto University.
Figure 7: Examples of NeKo outputs for OCR correction task in PleIAs/Post-OCR-Correction.
Figure 8: Examples of NeKo outputs for textual error correction (TEC) tasks in CoEdIT Raheja et al. (2023).
Figure 9: Examples of (a) Human recognition given different input modalities, including audio, text, and visual patterns; (b) generative inference and correction Marshall et al. (2015); Levinson and Evans (2010) to understand the recognition results.
Figure 10: We provide medical post-ASR recognition correction on the Medical-ASR-EN dataset (https://huggingface.co/datasets/jarvisx17/Medical-ASR-EN), where NeKo demonstrates the ability to (1) refine clinically related term errors and (2) correct grammar format.

### v2

@@
-Large Language Models with Task-Oriented Experts
+Yen-Ting Lin Zhehuai Chen Piotr Zelasko Zhen Wan Xuesong Yang Zih-Ching Chen Krishna C Puvvada Szu-Wei Fu  Ke Hu  Jun Wei Chiu  Jagadeesh Balam  Boris Ginsburg  Yu-Chiang Frank Wang  Chao-Han Huck Yang NVIDIA
 Abstract
@@
 We propose a Multi-Task Correction MoE, where we train the experts to become an “expert” of speech-to-text, language-to-text and vision-to-text datasets by learning to route each dataset’s tokens to its mapped expert.
-Experiments on the Open ASR Leaderboard show that we explore a new state-of-the-art performance by achieving an average relative $5.0$% WER reduction and substantial improvements in BLEU scores for speech and translation tasks.
-On zero-shot evaluation, NeKo outperforms GPT-3.5 and Claude-Opus with $15.5$% to $27.6$% relative WER reduction in the Hyporadise benchmark.
+Experiments on the Open ASR Leaderboard show that we explore a new state-of-the-art performance by achieving an average relative 5.05.0% WER reduction and substantial improvements in BLEU scores for speech and translation tasks.
+On zero-shot evaluation, NeKo outperforms GPT-3.5 and Claude-3.5 Sonnet with 15.515.5% to 27.627.6% relative WER reduction in the Hyporadise benchmark.
 NeKo performs competitively on grammar and post-OCR correction as a multi-task model.
-[Uncaptioned image]NeKo: Toward Post Recognition Generative Correction
-Large Language Models with Task-Oriented Experts
+[Uncaptioned image]NeKo: Cross-Modality Post-Recognition Error Correction with Tasks-Guided Mixture-of-Experts Language Model
 Yen-Ting Linthanks: Work done at NVIDIA research as an intern.
-Chao-Han Huck Yang  Zhehuai Chen Piotr Zelasko Xuesong Yang Zih-Ching Chen Krishna C Puvvada Szu-Wei Fu  Ke Hu  Jun Wei Chiu  Jagadeesh Balam  Boris Ginsburg  Yu-Chiang Frank Wang NVIDIA corresponding authors: ytl@ieee.org, hucky@nvidia.com
+Zhehuai Chen  Piotr Zelasko  Zhen Wan  Xuesong Yang Zih-Ching Chen  Krishna C Puvvada  Szu-Wei Fu   Ke Hu   Jun Wei Chiu   Jagadeesh Balam   Boris Ginsburg   Yu-Chiang Frank Wang   Chao-Han Huck Yang NVIDIA corresponding authors: ytl@ieee.org, hucky@nvidia.com
 1 Introduction
@@
 NeKo discover new state-of-the-art results in (iv) zero-shot ASR correction and performs competitively as a general-purpose (v) multi-task corrector.
-Human recognition (Biederman, 1987; Juang and Furui, 2000; Kanwisher et al., 1996) capabilities span multiple modalities, including speech recognition, visual patterns, and extensions to semantic and textual interpretations.
+Human recognition capabilities span multiple modalities, including speech recognition, visual patterns, and extensions to semantic and textual interpretations.
 These faculties, however, are not infallible and often incorporate mis-recognition errors.
@@
 In neuroscience Zatorre and Gandour (2008), the inferior temporal gyrus and the temporal lobe are not confined to rudimentary perception but are also integral to the post-recognition processes that facilitate semantic understanding of language (Levinson and Evans, 2010), speech (Marshall et al., 2015), and visual patterns (Vink et al., 2020).
-This form of “post-recognition correction,” exemplified by the application of language modeling (LM) to initial recognition outputs, has been introduced to the field for both acoustic (automatic speech recognition, ASR) and visual (optical character recognition, OCR) modalities since the early explorations (Jelinek, 1976; Dixon and Silverman, 1975) of learning algorithms in 1970s.
-The most prevalent approaches to utilizing LMs for post-recognition boosting are predominantly ranking or retrieval-based.
-In these setups, the LM is tasked with ranking and scoring (Ljolje et al., 1999) the top $n$-best hypotheses generated by the first-pass recognition system.
-This process often incorporates “discriminative modeling” algorithms (Sukkar and Lee, 1996; Mangu et al., 2000) and representation embeddings, such as BERT (Salazar et al., 2020; Kenton and Toutanova, 2019), to minimize the word error rate (WER) during training (Prabhavalkar et al., 2018; Mangu et al., 2000).
+This form of “post-recognition correction,” exemplified by the application of language modeling (LM) to initial recognition outputs, has been introduced to the field for both acoustic (automatic speech recognition, ASR) and visual (optical character recognition, OCR) modalities.
 With the LMs scaling up to LLMs (Brown et al., 2020), recent efforts (Chan et al., 2023; Yang et al., 2023; CHEN et al., 2023; Hu et al., 2024a) have focused on exploring a “generative modeling” for post-recognition correction.
@@
 Meanwhile, these GER solutions heavily depends on domain-specific fine-tuning processes (Chen et al., 2024a) that utilize parameter-efficient components, which often suffers a performance degradation from a lack of generalizability across different datasets, domains, and tasks.
-In other words, how to design and further openly provide a “general-purpose post (every) recognition correction model” is still one undiscovered and crucial topic within the research community.
-On the other hand, directly fine-tuning LLMs on a mixture of diverse error correction datasets can lead to suboptimal performance (CHEN et al., 2023; Lange et al., 2022) due to differences in input modalities, output formats, error types, and domain characteristics.
-For example, ASR errors stem from phonetic similarities or acoustic ambiguities, while OCR errors involve visual or character-level confusions.
-Additionally, error distributions can vary widely across datasets, even within the same task.
 To characterize “model generalization,” mixture-of-experts (MoE) (Jiang et al., 2024a) has emerged as a promising approach for multi-task learning, consisting of of a set of expert networks and a gating network that learns to route the input to the most appropriate expert (Sukhbaatar et al., 2024).
 This enables MoE models to learn more specialized and fine-grained representations compared to monolithic models.
-However, most MoE models are designed for general-purpose language modeling(Dai et al., 2024), with experts not explicitly assigned to specific tasks, but rather learn to specialize in different aspects of the input space through data-driven training, expect for a recent vision work (Ye and Xu, 2023).
+However, most MoE models are designed for general-purpose language modeling(Dai et al., 2024), with experts not explicitly assigned to specific tasks, but rather learn to specialize in different aspects of the input space through data-driven training.
 Effectively leverage MoE for multi-task error correction, where the experts need to capture task-specific features while allowing knowledge sharing, remains an open question.
@@
 The key idea is to continuously pre-train MoE model on a mixture of error correction datasets, with each expert specializing in a specific domain.
-This task-oriented MoE fine-tuning approach enables the experts to capture task-specific features while allowing knowledge sharing through the router (Dai et al., 2024).
+This task-guided MoE fine-tuning approach enables the experts to capture task-specific features while allowing knowledge sharing through the router.
+We further pursue this direction by modeling MoE on error correction and highlight the effectiveness and robustness of MoEs in learning from a mixture of correction datasets.
 NeKo captures the nuances of each task, benefiting from shared knowledge across experts.
-Evaluated on tasks such as ASR, ST, OCR, and unseen textual error correction (TEC), NeKo consistently outperforms baseline models, including Claude-Opus and GPT-3.5.
+Evaluated on tasks such as ASR, ST, OCR, and unseen textual error correction (TEC), NeKo consistently outperforms baseline models, including Claude-3.5 Sonnet and GPT-3.5.
 It achieves state-of-the-art WER reduction on the Hyporadise benchmark and large-scale Open ASR Leaderboard (Srivastav et al., 2023).
@@
   1. 1.
-    We introduce NeKo, a multi-task error correction LLM that leverages task-oriented mixture-of-experts for diverse post-recognition correction tasks.
+    We introduce NeKo, a multi-task error correction LLM that leverages task-guided mixture-of-experts for diverse post-recognition correction tasks.
 To the best of our knowledge, this is the first work that explores the use of MoE for multi-task error correction.
@@
     NeKo has been studied under a new form of cross-modalities post-recognition correction evaluation, serving as strong open-source ASR, ST, OCR, and TEC baselines.
-Our results show that NeKo discovers new state-of-the-art performance in ASR as a multi-task error correction model.
+Our results show that NeKo discovers new state-of-the-art performance in ASR as a multi-task correction model.
   3. 3.
@@
   4. 4.
-    The NeKo models, newly created source datasets, and training processes are scheduled to open source under the CC BY-SA 4.0 license to support reproducibility and to encourage future research.
-2 Related Work
-Language Modeling and Generative Error Correction
-Neural correction LMs have been widely used for end-to-end (E2E) models for text error correction or normalization for both ASR (Jelinek, 1990; Irie et al., 2016; Zhang et al., 2019a, b; Guo et al., 2019) and OCR (Plamondon and Srihari, 2000; Sabir et al., 2017; Fogel et al., 2020).
-These models often use beam search to generate new estimates, and can usually handle text normalization and denormalization or spelling errors.
-To utilize the abundant textual data to improve ASR, neural LMs are integrated to E2E models using shallow fusion or rescoring over acoustic model’ likelihood confidence (Kannan et al., 2018; Salazar et al., 2020; Yang et al., 2021).
-In addition to using only textual hypotheses for correction, deliberation models (Hu et al., 2020; Wang et al., 2022; Hu et al., 2023a) utilizes both audio and text for generating new ASR hypotheses.
-While the aforementioned methods have drawn much attention in the past, recent advances focus on using pretrained textual LLMs to benefit E2E ASR (Song et al., 2023; Hu et al., 2023c), or similar to deliberation, employing both speech and text for speech understanding by prompting (Gong et al., 2023) and joint speech and language foundation models (SLMs) (Wang et al., 2023).
-On the other hand, text-to-text based GEC (CHEN et al., 2023; Yang et al., 2023; Radhakrishnan et al., 2023; Chan et al., 2023) have been recently introduced to ASR, ST, image captioning, and video summarization to set up superior performance  [^1] GEC have shown better results compared to a similar scale of (i) multi-modal injection model and (ii) re-ranking based model, where we will leave these studies in the future upon having one multi-task correction LM as NeKo presented here.
-Please also.
-These GEC approaches take the text output as an input to large language models (LLMs), enabling them to perform zero-shot typo correction by generatively refining the final recognition results (CHEN et al., 2023; Yang et al., 2023; Radhakrishnan et al., 2023; Chan et al., 2023).
-In this work, we aim to design a text-to-text model as a first step for providing open-source post-recognition correction to the communities.
-Mixture of Experts
-Mixture-of-experts (MoE) (Shazeer et al., 2017) is a machine learning concept that employs multiple expert layers, each of which specializes in solving a specific subtask.
-The experts then work together to solve the entire task at hand.
-Recently, MoE has been widely applied to large-scale distributed Deep Learning models by using a cross-GPU layer that exchanges hidden features from different GPUs (Lepikhin et al., 2021; Fedus et al., 2022).
-The MoE approach is differentiated from existing scale-up approaches for DNNs, such as increasing the depth or width of DNNs, in terms of its high cost-efficiency.
-Specifically, adding more model parameters (experts) in MoE layers does not increase the computational cost per token at inference time.
-Thus, MoE has been studied for scaling the models to trillion-size parameters in NLP (Fedus et al., 2022).
-In speech processing, Gaur et al.
-(2021); Hu et al.
-(2023b); You et al.
-(2021) primarily showed benefits of MoE on speech recognition tasks.
-In this work, we further pursue this direction by modeling MoE on error correction and highlight the effectiveness and robustness of MoEs in learning from a mixture of correction datasets.
+    The NeKo models, newly created source datasets, and training processes are scheduled to open source under the CC BY-SA 4.0 license to support reproducibility in future research.
 2 Method
@@
 The number of experts $K$ used per token is a hyperparameter that controls the computational cost.
-3.2 Task-Oriented Expert Assignment
+2.2 Tasks-Guided Auxiliary Expert Assignment
 The key idea of NeKo is to assign each expert to a specific task during training.
@@
 $y=G(x)_{f(T_{i})}\cdot E_{f(T_{i})}(x)+G(x)_{\text{top1}}\cdot E_{\text{top1}}(x),$ (3)
-[…]
+where $\text{top1}=\operatorname*{arg\,max}_{j\neq f(T_{i})}G(x)_{j}$ is the index of the top-1 expert selected by the gating network, excluding the task-specific expert $f(T_{i})$.
 During inference, we do not assume knowledge of the specific task an input token belongs to.
@@
 $\mathcal{L}=-\sum_{i=1}^{m}\sum_{(x,y)\in D_{i}}\log p(y|x,T_{i}),$ (4)
-where $x$ is the input sequence (e.g., ASR hypotheses, OCR output), $y$ is the target sequence (e.g., ground-truth transcription, corrected text), and $p(y|x,T_{i})$ is the probability of the target sequence given the input sequence and the task prompt (Figure 3.) By jointly training on multiple error correction datasets with task-oriented expert assignment, NeKo learns to capture task-specific features while allowing for knowledge sharing across tasks through the shared gating network and other model components.
+where $x$ is the input sequence (e.g., ASR hypotheses, OCR output), $y$ is the target sequence (e.g., ground-truth transcription, corrected text), and $p(y|x,T_{i})$ is the probability of the target sequence given the input sequence and the task prompt (Figure 3.) By jointly training on multiple error correction datasets with task-guided expert assignment, NeKo learns to capture task-specific features while allowing for knowledge sharing across tasks through the shared gating network and other model components.
 3 Experiments
@@
 ST and MT
-For the speech translation error correction task, we use the subset of the HypoTranslate dataset (Hu et al., 2024b) for training and evaluation.
-This dataset includes translation results from FLEURS (Conneau et al., 2022), CoVoST-2 (Wang et al., 2020), and MuST-C (Di Gangi et al., 2019), covering a range of languages such as Spanish, French, Italian, Japanese, Portuguese, Chinese, and Persian.
-As an extra zero-shot textual correction setup, we evaluate NeKo on machine translation (MT) of WMT’20 for Japanese and Chinese (Barrault et al., 2020a).
-We use the BLEU score (Papineni et al., 2002) as the evaluation metric for ST (with training and test) and MT (zero-shot).
+For the translation error correction task, we use the subset of the HypoTranslate dataset (Hu et al., 2024b) for training and evaluation.
+This dataset includes translation from FLEURS (Conneau et al., 2022), CoVoST-2 (Wang et al., 2020), and MuST-C (Di Gangi et al., 2019), covering a range of languages such as Spanish, French, Italian, Japanese, Portuguese, Chinese, and Persian.
 OCR
-For the optical character recognition (OCR) error correction task, we use the English portion of the Post-OCR Correction dataset (PleIAs, 2023), which contains newspaper texts from Chronicling America.
-The dataset includes original texts with varying numbers of OCR mistakes and their corresponding corrected versions.
-To evaluate our model, we take the first 1,000 characters of both the input text with OCR errors and the ground-truth corrected text.
-We use the WER as the evaluation metric.
+For the optical character recognition (OCR) error correction task, we use the en-us portion of the OCR dataset (PleIAs, 2023), which contains newspaper texts from Chronicling America.
 TEC
-For the textual error correction (TEC) task, we use a subset of the CoEdIT dataset (Raheja et al., 2023), which contains a diverse collection of 82K task-specific instructions for text editing.
-We select two editing tasks from CoEdIT that align with our error correction objectives: grammar correction and coherence improvement.
-These tasks focus on correcting grammatical errors and improving the overall coherence of the text, making them suitable for evaluating the effectiveness of our model in handling TEC-related editing instructions.
-We use the word error rate as the evaluation metric.
+For the textual error correction (TEC) task, we use a subset of the CoEdIT dataset (Raheja et al., 2023) from Grammarly, which contains 82K task-specific instructions for text editing.
 3.2 Task-Specific Recognition Systems and Baselines
 ASR
-We compare against state-of-the-art ASR models, Distil-Whisper-V2-Large(Gandhi et al., 2023), Whisper-V2-Large, Whisper-V3-Large(Radford et al., 2022), Canary(NVIDIA, 2024) without applying GEC method.
-A end-to-end ASR-LLM, SALM (Chen et al., 2024b), is also compared.
+We compare against state-of-the-art ASR models, Whisper-V2-Large (Radford et al., 2022), Canary (NVIDIA, 2024) without applying GEC method.
+End-to-end ASR-LLM, SALM (Chen et al., 2024b), Qwen2-audio, and Gemini-2-Flash have been also compared.
 For all Cascaded ASR+GEC Methods, the task-specific system is the Canary model.
 This model transcribes the speech data and generate 5-best hypotheses for each utterance using temperature-based sampling (Ackley et al., 1985) with $p=0.3$.
-This allows us to capture a diverse set of potential transcriptions for each utterance, which can then be fed into our error correction model.
+This allows us to capture a diverse set of potential transcriptions for each utterance, which can be fed into our error correction model.
 ST and MT
@@
 We compare our proposed method against two baselines: (1) the input text without any correction (denoted as Baseline) and (2) a Mixtral 8x7B model fine-tuned only on the respective dataset for each task (denoted as Mistral 8x7B Direct Finetune).
-This allows us to assess the effectiveness of our task-oriented expert assignment approach in handling OCR and TEC errors, as well as its ability to leverage knowledge from multiple tasks to improve performance on individual tasks compared to direct fine-tuning on a single dataset.
+This allows us to assess the effectiveness of our task-guided expert assignment approach in handling OCR and TEC errors, as its ability to leverage knowledge from multiple tasks to improve performance on individual tasks compared to direct fine-tuning on a single dataset.
 Table 1: Cross-domain ASR correction results in zero-shot and few-shot settings on the Hyporadise benchmark (CHEN et al., 2023).
-We compare NeKo against GPT-3.5 Turbo and Claude-Opus in 0- and 5-shot settings.
+We compare NeKo against GPT-4 Turbo and Claude-3.5 Sonnet in 0- and 5-shot settings.
 The baseline represents the WER of task-specific model Whisper-Large.
@@
 (2023) (N-best and Compositional) provide an upper bound for the correction performance.
-Domain Test Set Baseline GPT-3.5 Turbo Claude-Opus 0-shot w/ NeKo Oracle
+Domain Test Set Baseline GPT-3.5 Turbo Claude-3.5 Sonnet 0-shot w/ NeKo Oracle
 Shift 0-shot 5-shot 0-shot 5-shot NeKo-FFT NeKo-BTX NeKo-MoE N-best Comp.
-Specific Scenario WSJ-dev93 9.0 $8.5_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $7.7_{{\color[rgb]{0,.5,.5}-14.4\%}}$ $8.2_{{\color[rgb]{0,.5,.5}-8.9\%}}$ $7.4_{{\color[rgb]{0,.5,.5}-17.8\%}}$ $8.6_{{\color[rgb]{0,.5,.5}-4.4\%}}$ $7.5_{{\color[rgb]{0,.5,.5}-16.7\%}}$ $\textbf{6.8}_{{\color[rgb]{0,.5,.5}-24.4\%}}$ 6.5 5.3
+Specific
+Scenario
+WSJ-dev93 9.0 $8.5_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $7.7_{{\color[rgb]{0,.5,.5}-14.4\%}}$ $8.2_{{\color[rgb]{0,.5,.5}-8.9\%}}$ $7.4_{{\color[rgb]{0,.5,.5}-17.8\%}}$ $8.6_{{\color[rgb]{0,.5,.5}-4.4\%}}$ $7.5_{{\color[rgb]{0,.5,.5}-16.7\%}}$ $\textbf{6.8}_{{\color[rgb]{0,.5,.5}-24.4\%}}$ 6.5 5.3
 WSJ-eval92 7.6 $7.3_{{\color[rgb]{0,.5,.5}-3.9\%}}$ $6.6_{{\color[rgb]{0,.5,.5}-13.2\%}}$ $7.0_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $6.3_{{\color[rgb]{0,.5,.5}-17.1\%}}$ $7.4_{{\color[rgb]{0,.5,.5}-2.6\%}}$ $6.4_{{\color[rgb]{0,.5,.5}-15.8\%}}$ $\textbf{5.8}_{{\color[rgb]{0,.5,.5}-23.7\%}}$ 5.5 4.7
 ATIS 5.8 $5.5_{{\color[rgb]{0,.5,.5}-5.2\%}}$ $5.0_{{\color[rgb]{0,.5,.5}-13.8\%}}$ $5.2_{{\color[rgb]{0,.5,.5}-10.3\%}}$ $4.7_{{\color[rgb]{0,.5,.5}-19.0\%}}$ $5.6_{{\color[rgb]{0,.5,.5}-3.4\%}}$ $4.8_{{\color[rgb]{0,.5,.5}-17.2\%}}$ $\textbf{4.2}_{{\color[rgb]{0,.5,.5}-27.6\%}}$ 3.5 2.4
-Common Noise CHiME4-bus 18.8 $17.6_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $16.2_{{\color[rgb]{0,.5,.5}-13.8\%}}$ $17.1_{{\color[rgb]{0,.5,.5}-9.0\%}}$ $15.7_{{\color[rgb]{0,.5,.5}-16.5\%}}$ $17.7_{{\color[rgb]{0,.5,.5}-5.9\%}}$ $15.9_{{\color[rgb]{0,.5,.5}-15.4\%}}$ $\textbf{14.5}_{{\color[rgb]{0,.5,.5}-22.9\%}}$ 16.8 10.7
+Common
+Noise
+CHiME4-bus 18.8 $17.6_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $16.2_{{\color[rgb]{0,.5,.5}-13.8\%}}$ $17.1_{{\color[rgb]{0,.5,.5}-9.0\%}}$ $15.7_{{\color[rgb]{0,.5,.5}-16.5\%}}$ $17.7_{{\color[rgb]{0,.5,.5}-5.9\%}}$ $15.9_{{\color[rgb]{0,.5,.5}-15.4\%}}$ $\textbf{14.5}_{{\color[rgb]{0,.5,.5}-22.9\%}}$ 16.8 10.7
 CHiME4-caf 16.1 $14.7_{{\color[rgb]{0,.5,.5}-8.7\%}}$ $13.7_{{\color[rgb]{0,.5,.5}-14.9\%}}$ $14.2_{{\color[rgb]{0,.5,.5}-11.8\%}}$ $13.2_{{\color[rgb]{0,.5,.5}-18.0\%}}$ $14.8_{{\color[rgb]{0,.5,.5}-8.1\%}}$ $13.4_{{\color[rgb]{0,.5,.5}-16.8\%}}$ $\textbf{12.2}_{{\color[rgb]{0,.5,.5}-24.2\%}}$ 13.3 9.1
@@
 CHiME4-str 11.4 $10.9_{{\color[rgb]{0,.5,.5}-4.4\%}}$ $9.7_{{\color[rgb]{0,.5,.5}-14.9\%}}$ $10.5_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $9.3_{{\color[rgb]{0,.5,.5}-18.4\%}}$ $11.0_{{\color[rgb]{0,.5,.5}-3.5\%}}$ $9.4_{{\color[rgb]{0,.5,.5}-17.5\%}}$ $\textbf{8.5}_{{\color[rgb]{0,.5,.5}-25.4\%}}$ 9.0 6.0
-Speaker Accent MCV-af 25.3 $24.9_{{\color[rgb]{0,.5,.5}-1.6\%}}$ $23.6_{{\color[rgb]{0,.5,.5}-6.7\%}}$ $24.4_{{\color[rgb]{0,.5,.5}-3.6\%}}$ $23.0_{{\color[rgb]{0,.5,.5}-9.1\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-1.2\%}}$ $23.3_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $\textbf{21.0}_{{\color[rgb]{0,.5,.5}-17.0\%}}$ 23.6 21.7
+Speaker
+Accent
+MCV-af 25.3 $24.9_{{\color[rgb]{0,.5,.5}-1.6\%}}$ $23.6_{{\color[rgb]{0,.5,.5}-6.7\%}}$ $24.4_{{\color[rgb]{0,.5,.5}-3.6\%}}$ $23.0_{{\color[rgb]{0,.5,.5}-9.1\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-1.2\%}}$ $23.3_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $\textbf{21.0}_{{\color[rgb]{0,.5,.5}-17.0\%}}$ 23.6 21.7
 MCV-au 25.8 $25.1_{{\color[rgb]{0,.5,.5}-2.7\%}}$ $24.0_{{\color[rgb]{0,.5,.5}-7.0\%}}$ $24.6_{{\color[rgb]{0,.5,.5}-4.7\%}}$ $23.4_{{\color[rgb]{0,.5,.5}-9.3\%}}$ $25.2_{{\color[rgb]{0,.5,.5}-2.3\%}}$ $23.7_{{\color[rgb]{0,.5,.5}-8.1\%}}$ $\textbf{21.4}_{{\color[rgb]{0,.5,.5}-17.1\%}}$ 24.9 21.8
-MCV-in 28.6 $27.6_{{\color[rgb]{0,.5,.5}-3.5\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-12.6\%}}$ $27.0_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $24.3_{{\color[rgb]{0,.5,.5}-15.0\%}}$ $27.8_{{\color[rgb]{0,.5,.5}-2.8\%}}$ $24.6_{{\color[rgb]{0,.5,.5}-14.0\%}}$ $\textbf{22.2}_{{\color[rgb]{0,.5,.5}-22.4\%}}$ 27.1 22.6
-MCV-sg 26.4 $26.5_{{\color[rgb]{.5,.5,.5}+0.4\%}}$ $25.1_{{\color[rgb]{0,.5,.5}-4.9\%}}$ $25.9_{{\color[rgb]{0,.5,.5}-1.9\%}}$ $24.5_{{\color[rgb]{0,.5,.5}-7.2\%}}$ $26.6_{{\color[rgb]{.5,.5,.5}+0.8\%}}$ $24.7_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $\textbf{22.3}_{{\color[rgb]{0,.5,.5}-15.5\%}}$ 25.5 22.2
+[…]
+[…]
 Table 2: ASR correction results on the Open ASR Leaderboard.
@@
 NeKo establishes a new state-of-the-art performance on the leaderboard, outperforming both end-to-end ASR methods and cascaded ASR+GEC approaches.
-Lower WER indicates better performance.
-Model AMI Earnings22 Gigaspeech LS Clean LS Other SPGI Tedlium Voxp. MCV9 Average
-End-to-end ASR Methods
-Distil-Whisper-V2-Large (Gandhi et al., 2023) 14.65 12.12 10.31 2.95 6.39 3.28 4.30 8.22 12.60 8.31
-Whisper-V2-Large (Radford et al., 2022) 16.82 12.02 10.57 2.56 5.16 3.77 4.01 7.50 10.11 8.06
-Whisper-V3-Large 16.01 11.30 10.02 2.03 3.91 2.95 3.90 9.52 9.67 7.70
-Canary (NVIDIA, 2024) 14.00 12.25 10.19 1.49 2.49 2.06 3.58 5.81 7.75 6.67
-Bestow Speech LM (Chen et al., 2024c) 12.58 12.86 10.06 1.64 3.07 2.11 3.41 5.84 6.97 6.50
-Cascaded ASR+GEC Methods
-Gemma 2B (Team et al., 2024) FFT 13.20 12.30 10.40 1.60 2.60 2.20 3.70 6.00 7.50 6.61
-Gemma 8x2B FFT 13.10 12.20 10.30 1.50 2.50 2.10 3.60 5.90 7.40 6.51
-NeKo (Ours) Gemma 8x2B 13.00 12.10 10.20 1.40 2.40 2.00 3.50 5.80 7.30 6.41
-Mistral 7B (Jiang et al., 2023) FFT 13.07 11.87 10.09 1.48 2.46 2.04 3.55 5.75 7.29 6.40
-Mixtral 8x7B (Jiang et al., 2024b) FFT 12.91 12.19 10.34 1.54 2.55 2.12 3.64 5.89 7.43 6.51
-Mixtral 8x7B Lora 12.96 12.24 10.38 1.55 2.56 2.13 3.66 5.92 7.47 6.60
-Mistral 8x7B Branch-Train-Mix (Sukhbaatar et al., 2024) 13.13 11.93 10.14 1.49 2.47 2.05 3.57 5.78 7.33 6.43
-NeKo (Ours) Mixtral 8x7B 12.55 11.82 10.02 1.49 2.47 2.05 3.52 5.76 7.25 6.34
-NeKo (Ours) Mixtral 8x22B 12.61 11.93 10.15 1.52 2.51 2.09 3.58 5.82 7.33 6.40
+We report the actual tuning parameter in parentheses (.) and the sum of the frozen Whisper results in front.
+Model Inference Para. Avg. $\downarrow$ AMI Earnings22 Gigaspeech LS Clean LS Other SPGI Tedlium Voxp. MCV9
+ASR or SpeechLMs: End-to-end Voice Understanding Models
+Distil-Whisper-V2-L (Gandhi et al., 2023) 0.75B 8.31 14.65 12.12 10.31 2.95 6.39 3.28 4.30 8.22 12.60
+Whisper-V2-L (Radford et al., 2022) 1.5B 8.06 16.82 12.02 10.57 2.56 5.16 3.77 4.01 7.50 10.11
+Canary (NVIDIA, 2024) 2B 6.67 14.00 12.25 10.19 1.49 2.49 2.06 3.58 5.81 7.75
+Bestow Speech LM (Chen et al., 2024c) 1.8B 6.50 12.58 12.86 10.06 1.64 3.07 2.11 3.41 5.84 6.97
+Qwen2-Audio (Chu et al., 2024) 8B 7.43 - - - 1.6 3.6 - - - -
+Gemini-2.0-Flash - 8.56 - - - - - - - - -
+ASR+LLM: Frozen Whisper-v2-L (1.5B) + Voice Correction LMs
++ Gemma 2B (Team et al., 2024) FFT 3.5B (2B) 6.61 13.20 12.30 10.40 1.60 2.60 2.20 3.70 6.00 7.50
++ Gemma 8x2B FFT 3.5B (2B) 6.51 13.10 12.20 10.30 1.50 2.50 2.10 3.60 5.90 7.40
+NeKo (Ours) Gemma 8x2B 3.5B (2B) 6.41 13.00 12.10 10.20 1.40 2.40 2.00 3.50 5.80 7.30
+NeKo (Ours) Qwen1.5-MoE 4.2B (2.7B) 5.90 12.60 11.82 9.95 1.30 2.32 1.94 3.20 5.80 7.30
++ Mistral 7B (Jiang et al., 2023) FFT 8.5B (7B) 6.40 13.07 11.87 10.09 1.48 2.46 2.04 3.55 5.75 7.29
++ Mixtral 8x7B (Jiang et al., 2024b) FFT 8.5B (7B) 6.51 12.91 12.19 10.34 1.54 2.55 2.12 3.64 5.89 7.43
++ Mixtral 8x7B Lora 8.5B (7B) 6.60 12.96 12.24 10.38 1.55 2.56 2.13 3.66 5.92 7.47
++ Mistral 8x7B BTM (Sukhbaatar et al., 2024) 8.5B (7B) 6.43 13.13 11.93 10.14 1.49 2.47 2.05 3.57 5.78 7.33
+NeKo (Ours) Mixtral 8x7B 8.5B (7B) 6.34 12.55 11.82 10.02 1.49 2.47 2.05 3.52 5.76 7.25
+NeKo (Ours) Mixtral 8x22B 23.5B (22B) 6.40 12.61 11.93 10.15 1.52 2.51 2.09 3.58 5.82 7.33
 3.3 Post-recognition LLMs Setup
@@
 For MoE models, we fine-tune Gemma 8x2B [^1] We made an up-cycled (Komatsuzaki et al., 2023) Gemma 8x2B MoE setup extended from single Gemma-2B (Team et al., 2024).
-and Mixtral 8x7B without applying our task-oriented expert assignment.
+and Mixtral 8x7B without applying our task-guided expert assignment.
 We explore the Branch-Train-Mix approach (Sukhbaatar et al., 2024), which involves branching from the Mistral 7B model to an 8x7B MoE model as one competing setup.
@@
 We use a global batch size of 2 million tokens and apply sample packing (Raffel et al., 2020) to maximize the GPU utilization.
-We fine-tune the model for 3 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $41e-$ 4 and a weight decay of $0.01$.
-We use a cosine learning rate scheduler with a warmup ratio of $0.1$ and a gradient clipping threshold of $1.0$.
-For the expert-dataset mapping, we randomly assign each dataset to one of the 8 experts in the Mixtral model.
-This random assignment serves as a strong baseline and allows us to focus on the effectiveness of the task-oriented expert assignment approach.
-We leave the exploration of more advanced expert assignment strategies for future work.
-To efficiently train the large-scale model, we leverage DeepSpeed Zero (Rajbhandari et al., 2020) for memory optimization and Hugging Face Transformers (Wolf et al., 2020) for model implementation.
 […]
@@
 GenTranslate-V2 27.0 44.3 26.4 27.8 44.5 36.1 34.4 20.8 29.7 43.5 31.3 33.2 28.3 16.9 26.1
-NeKo-FT 26.9 44.2 26.3 27.7 44.4 36.0 34.3 20.7 29.6 43.4 31.2 33.1 28.2 16.8 26.0
-NeKo-BTX 27.2 44.5 26.7 28.0 44.7 36.3 34.6 21.0 29.9 43.8 31.6 33.4 28.5 17.1 26.3
-NeKo-MoE 28.5 46.2 28.0 30.1 46.3 38.7 36.3 23.4 32.6 46.5 34.2 37.2 32.8 21.5 30.5
-4.4 Post-recognition Generative Correction Results
+NeKo-Gemma-2B-FT 26.9 44.2 26.3 27.7 44.4 36.0 34.3 20.7 29.6 43.4 31.2 33.1 28.2 16.8 26.0
+NeKo-Gemma-8x2B–BTX 27.2 44.5 26.7 28.0 44.7 36.3 34.6 21.0 29.9 43.8 31.6 33.4 28.5 17.1 26.3
+NeKo-Gemma-8x2B-MoE 28.5 46.2 28.0 30.1 46.3 38.7 36.3 23.4 32.6 46.5 34.2 37.2 32.8 21.5 30.5
+3.4 Post-recognition Correction Results
 ASR
-We first evaluate the zero-shot ability of NeKo on unseen domain compared to two general-purpose LLMs, including GPT-3.5 Turbo and Claude-Opus [^3] Claude-Opus is one of the high-performance, general-purpose LLMs evaluated in the LLM evaluation (Zheng et al., 2024) and performs competitively compared to GPT-4-Turbo and Gemini 1.5-Pro, offering unlimited API access..
-With a task-specific recognition baseline of Whisper-V2-Large (third column) in Table 1, NeKo-MoE shows the best zero-shot ability with a relative 22.3% average WER reduction.
-GPT-3.5 Turbo and Claude-Opus have relative 4.3% and 7.3% of zero-shot improvements, where NeKo consistently outperform their 5-shot ASR correction.
+We first evaluate the zero-shot ability of NeKo on unseen domain compared to two general-purpose LLMs, including GPT-3.5 Turbo and Claude-3.5 Sonnet.
+With a task-specific recognition baseline of Whisper-V2-Large (third column) in Table 1, NeKo-MoE (i.e., Qwen1.5-MoE or Mixtral) shows the best zero-shot ability with a relative 22.3% average WER reduction.
+GPT-3.5 Turbo and Claude-3.5 Sonnet have relative 4.3% and 7.3% of zero-shot improvements, where NeKo consistently outperform their 5-shot ASR correction.
 Table 2 shows the WER scores on individual datasets and average performance on the Open ASR Leaderboard.
@@
 While, Earnings22 shows a slight performance drop possibly due to the reduced representation in the batch.
-Compared to other leading models on the leaderboard, NeKo establishes a new state-of-the-art, outperforming speech-only foundational models like Whisper and Canary and end-to-end ASR-LLM like SALM  [^4] We conduct one competitive end-to-end, multi-modal baseline based on Bestow (Chen et al., 2024c) with joint audio and text injection under cross-attention, which shows improved performance on the Open ASR benchmark.
-(Chen et al., 2024b) across most datasets.
+Compared to other leading models on the leaderboard, NeKo establishes a new state-of-the-art, outperforming speech-only foundational models like Whisper and Canary and end-to-end ASR-LLM like SALM (Chen et al., 2024b) across most datasets.
 On the AMI dataset, NeKo achieves a WER of 12.58%, significantly lower than Whisper’s 16.82%.
@@
 Moreover, NeKo outperforms other correction baselines, including the state-of-the-art GenTranslate model.
-On the FLEURS dataset, NeKo obtains an average BLEU score of 36.3, surpassing GenTranslate by 2.3 points.
-For CoVoST-2, NeKo achieves an average BLEU score of 34.2, outperforming GenTranslate by 3.2 points.
-On MuST-C, NeKo improves over GenTranslate by 3.2 BLEU points, reaching an average score of 30.5.
-Table 4: Machine translation BLEU scores on the WMT’20 Japanese (Ja) and Chinese (Zh) test sets (Barrault et al., 2020b).
+Table 4: Machine translation BLEU scores on the WMT’20 Japanese (Ja) and Chinese (Zh) test sets (Barrault et al., 2020a).
 NeKo is evaluated in a zero-shot setting, while other models are fine-tuned on the respective language pairs.
@@
 GenTranslate (fine-tuned) 21.4 30.7 26.1
-NeKo-MoE (zero-shot) 18.1 27.6 22.9
+NeKo-Gemma-MoE (zero-shot) 18.1 27.6 22.9
 To further assess the generalization ability of NeKo , we evaluate it on the WMT’20 machine translation benchmark for Japanese and Chinese in a zero-shot setting.
 As shown in Table 4, NeKo achieves competitive performance compared to fine-tuned MT models, obtaining an average BLEU score of 22.9.
-This result highlights the potential of NeKo to handle unseen translation tasks by leveraging the knowledge learned from speech translation datasets.
+This result highlights the potential of NeKo to handle unseen translation tasks by leveraging the knowledge learned from pre-training.
 OCR and TEC
@@
 This represents a significant improvement over the baseline and demonstrates the model’s ability to correct OCR errors effectively.
-Compared to the Mixtral-MoE model fine-tuned directly on the OCR dataset, NeKo obtains a 1.02% lower WER, highlighting the benefit of the task-oriented expert assignment approach.
+Compared to the Mixtral-MoE model fine-tuned directly on the OCR dataset, NeKo obtains a 1.02% lower WER, highlighting the benefit of the task-guided expert assignment approach.
 In the TEC task, NeKo showcases its versatility by improving the performance on both grammar correction and coherence improvement subtasks.
@@
 Table 5: WER comparison of NeKo against the baseline and a directly fine-tuned Mixtral-MoE model (8x7B) on grammar correction and coherence improvement tasks from the CoEdIT dataset (Raheja et al., 2023), and the OCR task using the PleIAs/Post-OCR-Correction dataset (PleIAs, 2023).
-Lower WER indicates better performance.
-Task Grammar Correction Coherence Improv. OCR
-Baseline 31.41 13.48 71.03
-Mixtral-MoE-FFT 10.73   9.25 15.45
-NeKo-MoE   9.42   9.71 14.43
-Table 6: WER comparison of NeKo against GPT-3.5-Turbo, and Claude-Opus on the 5-shot IMDb typographical error correction dataset (Shah and de Melo, 2020).
-The baseline represents the WER between the corrupted text and the ground truth.
-Lower WER indicates better performance in correcting typographical errors.
-Model WER
-Baseline (Corrupt vs Ground Truth) 18.35%
-GPT-3.5-Turbo (5-shots) 12.72%
-Claude-Opus (5-shots) 8.18%
-NeKo-MoE (5-shots) 11.62%
-Emergent Unseen Task Zero-Shot Performance
-We investigate NeKo’s generalization capabilities to unseen tasks using an additional synthetic typographical error correction dataset (Shah and de Melo, 2020).
-This dataset is derived from the IMDb test split, featured low noise levels (3.75% character error rate) with corruption applied using algorithms proposed in (Shah and de Melo, 2020).
-Our evaluation focused on zero-shot and five-shot learning scenarios to assess the adaptability of various models without and with minimal task-specific training.
-In the zero-shot scenario, where models were prompted to switch from an ASR task to typo correction without additional training, the challenge proved significant.
-The models, including the advanced Claude-Opus, yielded WERs above 30%.
-The predictions were markedly irrelevant to the ground truth, highlighting the difficulty of adapting to typo correction without specific fine-tuning.
-This finding prompts further investigation into efficient and effective training techniques for generalizing model capabilities across diverse linguistic tasks.
-In the five-shot scenario, all models improved against the corrupted baseline with Claude-Opus performing best.
-Notably, NeKo outperformed GPT-3.5-Turbo, indicating some affinity towards this task.
+Task / WER $\downarrow$ Grammar Correction Coherence Improv. OCR
+Mixtral-MoE (frozen) 31.41 13.48 71.03
+GPT-3.5-turbo 17.43   12.25 39.45
+Mixtral-MoE-FFT 10.73   12.05 45.32
+NeKo-Mixtral-MoE   9.42   9.71 14.43
 4 Conclusion
-In this work, we proposed NeKo, a multi-task GER approach that leverages task-oriented MoEs to handle diverse tasks.
+In this work, we proposed NeKo, a multi-task GER approach that leverages task-guided MoEs to handle diverse tasks.
 NeKo assigns each expert to a specific dataset during training, enabling the experts to capture task-specific features while allowing knowledge sharing through the gating network.
-Our results show that task-oriented expert assignment is a promising approach for multi-task learning in error correction and other natural language processing tasks.
+Our results show that task-guided expert assignment is a promising approach for multi-task learning in error correction and other natural language processing tasks.
 By aligning experts with datasets, NeKo can effectively capture the nuances and specificities of each task while benefiting from the shared knowledge learned by the gating network and other model components.
 Future work includes exploring more advanced expert assignment strategies, such as dynamically assigning experts based on the input characteristics.
-Investigating the interpretability of the learned expert representations and routing decisions would be important next step upon the open NeKo models.
-Limitation
+Ethical Considerations
 We aim to provide a transparent and comprehensive understanding of the current scope of NeKo, and pave the way for future research to further improve the NeKo model.
@@
 (2021), to better evaluate the robustness and generalizability of the model.
-Ethical and Societal Considerations
+Societal Considerations
 The study does not extensively address the ethical and societal implications of deploying NeKo in real-world applications.
@@
 Future work should also consider optimizing the training process to minimize the environmental impact, promoting sustainable AI development practices.
-Task-Specific Fine-Tuning
-The NeKo model employs task-oriented MoE fine-tuning, where each expert is assigned to a specific dataset.
-This approach may lead to overfitting to the specific characteristics of the training datasets even though knowledge could be shared.
-As a result, the model’s performance might degrade when applied to new tasks or datasets that were not part of the training set, limiting its adaptability.
-Investigating more dynamic and adaptive fine-tuning strategies that can generalize better across unseen tasks and datasets would be beneficial.
-Techniques such as meta-learning or continual learning could be explored to enhance the model’s adaptability and robustness.
-Future Connections to In-Context and Auto-Agent Learning with NeKo
-Integrating in-context learning (ICL) with NeKo could enable the model to adapt to various error correction tasks by conditioning on input examples without requiring explicit fine-tuning.
-This approach is particularly beneficial in scenarios where obtaining large labeled datasets for fine-tuning is impractical.
-By leveraging ICL, NeKo could adapt to diverse error types and use in-context examples to correct errors specific to new domains or applications, thereby improving its generalizability to real-world data.
-Furthermore, ICL would allow the model to dynamically adjust its error correction strategies based on the input context, enhancing its robustness to varying error distributions.
 Appendix A Appendix
+Training Details
+We fine-tune the model for 3 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of 1e41e-4 and a weight decay of $0.01$.
+We use a cosine learning rate scheduler with a warmup ratio of $0.1$ and a gradient clipping threshold of $1.0$.
+For the expert-dataset mapping, we randomly assign each dataset to one of the 8 experts in the Mixtral model.
+This random assignment serves as a strong baseline and allows us to focus on the effectiveness of the task-guided expert assignment approach.
+We leave the exploration of more advanced expert assignment strategies for future work.
+To efficiently train the large-scale model, we leverage DeepSpeed Zero (Rajbhandari et al., 2020) for memory optimization and Hugging Face Transformers (Wolf et al., 2020) for model implementation.
+Translation Tasks.
+As an extra zero-shot textual correction setup, we evaluate NeKo on machine translation (MT) of WMT’20 for Japanese and Chinese (Barrault et al., 2020b).
+We use the BLEU score (Papineni et al., 2002) as the evaluation metric for ST (with training and test) and MT (zero-shot).
+Grammar Correction Tasks.
+These text error correction (TEC) tasks focus on correcting grammatical errors and improving the overall coherence of the text, making them suitable for evaluating the effectiveness of our model in handling TEC-related editing instructions.
+We use the word error rate as the evaluation metric.
+OCR Tasks.
+The dataset includes original texts with varying numbers of OCR mistakes and their corresponding corrected versions.
+To evaluate our model, we take the first 1,000 characters of both the input text with OCR errors and the ground-truth corrected text.
+We use the WER as the evaluation metric.
+Mixture of Experts Background
+Mixture-of-experts (MoE) (Shazeer et al., 2017) is a machine learning concept that employs multiple expert layers, each of which specializes in solving a specific subtask.
+The experts then work together to solve the entire task at hand.
+Recently, MoE has been widely applied to large-scale distributed Deep Learning models by using a cross-GPU layer that exchanges hidden features from different GPUs (Lepikhin et al., 2021; Fedus et al., 2022).
+The MoE approach is differentiated from existing scale-up approaches for DNNs, such as increasing the depth or width of DNNs, in terms of its high cost-efficiency.
+Specifically, adding more model parameters (experts) in MoE layers does not increase the computational cost per token at inference time.
+Thus, MoE has been studied for scaling the models to trillion-size parameters in NLP (Fedus et al., 2022).
 Prompt Format
@@
 Figure 5: Examples of NeKo outputs for speech translation correction task in FLEURS Conneau et al. (2022).
-Figure 6: Examples of NeKo outputs for machine translation correction task in WMT20 Barrault et al. (2020a).
+Figure 6: Examples of NeKo outputs for machine translation correction task in WMT20 Barrault et al. (2020b).
 Additional Discussion on Human Recognition from Speech and Text Inputs
@@
 (2021), requiring efforts from both speech, NLP, and cognitive science communities as one human recognition example shown in Figure 9.
-Task-Oriented Inference for Mixture of Expert Models
+task-guided Inference for Mixture of Expert Models
 During inference, the Neko-model utilizes top-2 expert routing, instead of just top-1. Our pilot studies showed that top-1 routing indeed led to worse performance due to limited knowledge sharing.
-Using more than two experts (e.g., top-3 or higher) diverged from the training setup and increased inference costs (ranging from $23.5$% to $75.5$%) without significant gain (i.e., a relative difference of less than $0.06$%).
+Using more than two experts (e.g., top-3 or higher) diverged from the training setup and increased inference costs (ranging from 23.523.5% to 75.575.5%) without significant gain (i.e., a relative difference of less than 0.060.06%).
 Future Model Maintenance Plan and ASR Community
@@
 In general, NeKo’s post-ASR correction improvements are consistent across datasets and first-pass models, suggesting the benefits generalize beyond model-specific (i.e., Canary, Whisper, or SeamlessM4T) ’s strengths as the initial medical term correction results shown in Figure 10.
+Emergent Unseen Task Zero-Shot Performance
+We investigate NeKo’s generalization capabilities to unseen tasks using an additional synthetic typographical error correction dataset (Shah and de Melo, 2020).
+This dataset is derived from the IMDb test split, featured low noise levels (3.75% character error rate) with corruption applied using algorithms proposed in (Shah and de Melo, 2020).
+Our evaluation focused on zero-shot and five-shot learning scenarios to assess the adaptability of various models without and with minimal task-specific training.
+In the zero-shot scenario, where models were prompted to switch from an ASR task to typo correction without additional training, the challenge proved significant.
+The models, including the advanced Claude-Opus, yielded WERs above 30%.
+The predictions were markedly irrelevant to the ground truth, highlighting the difficulty of adapting to typo correction without specific fine-tuning.
+This finding prompts further investigation into efficient and effective training techniques for generalizing model capabilities across diverse linguistic tasks.
+In the five-shot scenario, all models improved against the corrupted baseline with Claude-Opus performing best.
+Notably, NeKo outperformed GPT-3.5-Turbo, indicating some affinity towards this task.
+Task-Specific Fine-Tuning
+The NeKo model employs task-guided MoE fine-tuning, where each expert is assigned to a specific dataset.
+This approach may lead to overfitting to the specific characteristics of the training datasets even though knowledge could be shared.
+As a result, the model’s performance might degrade when applied to new tasks or datasets that were not part of the training set, limiting its adaptability.
+Investigating more dynamic and adaptive fine-tuning strategies that can generalize better across unseen tasks and datasets would be beneficial.
+Techniques such as meta-learning or continual learning could be explored to enhance the model’s adaptability and robustness.
+Future Connections to In-Context and Auto-Agent Learning with NeKo
+Integrating in-context learning (ICL) with NeKo could enable the model to adapt to various error correction tasks by conditioning on input examples without requiring explicit fine-tuning.
+This approach is particularly beneficial in scenarios where obtaining large labeled datasets for fine-tuning is impractical.
+By leveraging ICL, NeKo could adapt to diverse error types and use in-context examples to correct errors specific to new domains or applications, thereby improving its generalizability to real-world data.
+Furthermore, ICL would allow the model to dynamically adjust its error correction strategies based on the input context, enhancing its robustness to varying error distributions.
+Table 6: WER comparison of NeKo against GPT-3.5-Turbo, and Claude-Opus on the 5-shot IMDb typographical error correction dataset (Shah and de Melo, 2020).
+The baseline represents the WER between the corrupted text and the ground truth.
+Lower WER indicates better performance in correcting typographical errors.
+Model WER
+Baseline (Corrupt vs Ground Truth) 18.35%
+GPT-3.5-Turbo (5-shots) 12.72%
+Claude-3-Sonnet (5-shots) 12.18%
+Claude-3.5 Sonnet (5-shots) 8.18%
+NeKo-MoE (5-shots) 11.62%
 Acknowledgment
-The authors would like to acknowledge the initial feedback from Taejin Park, Boyi Li, and Yejin Choi at NVIDIA and the discussion on medical speech recognition evaluation with Zhen Wan and Chenhui Chu from Kyoto University.
+The authors would like to acknowledge the Taipei-1 high-performance computing support provided by the NVIDIA Taiwan Research and Development Center (TRDC) project, and the initial feedback from NVIDIA’s Eric Kang, Taejin Park, Boyi Li, and Yejin Choi.
 Figure 7: Examples of NeKo outputs for OCR correction task in PleIAs/Post-OCR-Correction.